# ToolSelf: Unifying Task Execution and Self-Reconfiguration via Tool-Driven Intrinsic Adaptation

### v1

@@ Abstract
Agentic systems powered by Large Language Models (LLMs) have demonstrated remarkable potential in tackling complex, long-horizon tasks. However, their efficacy is fundamentally constrained by static configurations governing agent behaviors, which are fixed prior to execution and fail to adapt to evolving task dynamics. Existing approaches, relying on manual orchestration or heuristic-based patches, often struggle with poor generalization and fragmented optimization. To transcend these limitations, we propose ToolSelf, a novel paradigm enabling tool-driven runtime self-reconfiguration. By abstracting configuration updates as a callable tool, ToolSelf unifies task execution and self-adjustment into a single action space, achieving a phase transition from external rules to intrinsic parameters. Agents can thereby autonomously update their sub-goals and context based on task progression, and correspondingly adapt their strategy and toolbox, transforming from passive executors into dual managers of both task and itself. We further devise Configuration-Aware Two-stage Training (CAT), combining rejection sampling fine-tuning with trajectory-level reinforcement learning to internalize this meta-capability. Extensive experiments across diverse benchmarks demonstrate that ToolSelf rivals specialized workflows while generalizing to novel tasks, achieving a 24.1% average performance gain and illuminating a path toward truly self-adaptive agents. [^1] Code is being organized.
@@ 1 Introduction
Agentic systems refer to large language model–based systems that go beyond passive text generation to autonomously plan, reason, and act in interactive environments through multi-step decision making (Kolt, 2025). By tightly coupling internal reasoning with external actions—such as tool invocation, environment interaction, and intermediate state updates—agentic systems enable models to tackle complex, long-horizon tasks that cannot be addressed through single-shot inference (Yao et al., 2022). This paradigm has become increasingly important as real-world applications, including deep research (Wu et al., 2025a), software engineering (Yang et al., 2024), and system control (Wang et al., 2024; Qin et al., 2025), demand sustained goal pursuit, adaptive planning, and iterative integration of information over time.
@@
Central to the behavior of an agentic system is its configuration, which defines the internal control variables governing how the agent reasons, acts, and interacts with its environment during execution. Configuration typically encompasses the agent’s execution strategy (e.g., persona or reasoning style) (Chen et al., 2023), the available toolbox of actions (Lu et al., 2025), and the task-relevant knowledge or context maintained across steps (Zhou et al., 2025). These elements implicitly shape the agent’s action space, information flow, and decision-making dynamics, and thus play a decisive role in long-horizon performance (Guo et al., 2024). However, most existing agentic systems rely on a static configuration that is fixed prior to execution and remains unchanged throughout the task. While adequate for short or deterministic questions, this rigidity becomes a fundamental bottleneck in long-horizon settings, where sub-goals evolve, information requirements shift, and tool needs change over time (Wölflein et al., 2025). Consequently, configuration is not merely an implementation detail, but a first-class component of agentic systems that fundamentally determines their autonomy, efficiency, and capacity for sustained reasoning.
@@
Prior work (Appendix 4) has explored two main directions to tackle these challenges. Multi-agent workflows (Hu et al., 2025; Fourney et al., 2024) rely on manually specified priors to decompose tasks in advance and orchestrate sub-agents with specialized configurations for collaborative execution. While effective in narrowly defined domains, the heavy reliance on human-designed decomposition, prompts, and orchestration logic severely limits their scalability, transferability, and generalization to unseen tasks. In contrast, single-agent extensions introduce external mechanisms for periodic configuration adjustment (Sun et al., 2023; Zhu et al., 2025) or interaction history compression (Wu et al., 2025b; Zhou et al., 2025). However, these solutions operate as fragmented patches targeting isolated symptoms rather than addressing the underlying cause. More critically, their triggering logic is typically governed by rigid heuristic thresholds—such as step counts or context length—which decouple configuration updates from task semantics and execution dynamics, thereby precluding end-to-end, data-driven optimization.
@@
To this end, we propose ToolSelf (Tool-driven Self-reconfiguration), a novel paradigm that unifies task execution and self-configuration into a single action space by abstracting configuration updates as a callable tool, thereby effecting a phase transition from externally imposed rules to intrinsically learned control parameters. As illustrated in Figure 1, ToolSelf departs from conventional static architectures by equipping the inference agent with a reconfiguration tool, enabling it to autonomously and dynamically update its sub-goals (Sun et al., 2023) and context in response to task progression and environmental feedback, while adaptively configuring appropriate execution strategies and toolsets for each task stage. Sub-goal updates provide macro-level guidance that mitigates blind exploration in exponentially growing search spaces (Zhou et al., 2023), while proactive context management ensures refined and efficient information processing (Wu et al., 2025b; Zhou et al., 2025). Together, these mechanisms elevate agents from passive executors to dual managers of both task execution and self-configuration.
To enable the data-driven acquisition of self-reconfiguration capabilities, we further devise Configuration-Aware Two-stage Training (CAT). Both the inference agent and the reconfiguration engine share a single base model, yielding a compact and end-to-end trainable architecture. CAT optimizes both modules through two stages: Stage I employs rejection sampling fine-tuning (RFT) (Li et al., 2025a) to initialize both modules using teacher-generated successful trajectories, as effective self-reconfiguration is difficult to explore from scratch. Stage II applies KTO-based reinforcement learning (Ethayarajh et al., 2024) to further optimize reconfiguration decisions. Since self-reconfiguration is a meta-capability across execution stages whose quality only manifests through downstream task performance, we adopt trajectory-level credit assignment—propagating task success/failure to all reconfiguration decisions within the trajectory—for which KTO’s binary feedback is naturally well aligned.
@@ 2.1 Preliminaries: Pre-Fixed Configuration Paradigm
Formally, the behavior policy of a ReAct-style (Yao et al., 2022) single-agent can be described as a conditional distribution:
@@ 2.3 Configuration-Aware Two-stage Training
To internalize the meta-capability of self-reconfiguration, we propose Configuration-Aware Two-stage Training (CAT) to jointly train both the inference agent $\pi_{i}$ and the reconfiguration engine $\mu$. CAT consists of Rejection Sampling Fine-tuning (RFT) (Li et al., 2025a) for cold-start initialization and Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024) for reinforcement learning refinement. Both $\pi_{i}$ and $\mu$ are fine-tuned via Low-Rank Adaptation (LoRA) (Hu et al., 2022) with independent adapters. Let $\theta_{0}$ denote the frozen base model parameters, and let $\Delta\theta_{\pi}$ and $\Delta\theta_{\mu}$ denote the LoRA parameters for the inference agent and reconfiguration engine, respectively. The effective parameters of the two modules are thus $\theta_{\pi}=\theta_{0}+\Delta\theta_{\pi}$ and $\theta_{\mu}=\theta_{0}+\Delta\theta_{\mu}$. Crucially, $\pi_{i}$ and $\mu$ share a single base model; LoRA adapters merely specialize it, yielding a compact and end-to-end trainable architecture.
Stage I: Cold-Start Initialization. To provide strong priors for configuration decisions, we bootstrap from teacher demonstrations (Li et al., 2025a). We generate trajectories using a teacher model on the task dataset and filter successful trajectories via rejection sampling. For task $Q^{(j)}$, let $\mathcal{T}_{\text{teacher}}^{(j)}$ denote the successful trajectory obtained through rejection sampling, and define a binary success label $y(\mathcal{T}_{\text{teacher}}^{(j)})\in\{0,1\}$, where 1 indicates successful task completion and 0 indicates failure. The filtered dataset is defined as:
@@ 2.4 Design Advantages
Autonomy. The inference agent retains full authority over macro-level task management. By formulating reconfiguration requests $r_{i}$—including proposed sub-goals $q_{i+1}^{\text{prop}}$—the agent itself drives task decomposition and progress tracking; the reconfiguration engine $\mu$ merely acts as a “compiler” translating these requests into concrete configurations. This contrasts with OWL (Hu et al., 2025) and OAgents (Zhu et al., 2025), which delegate planning to a separate high-level module, incurring additional coordination overhead.
Unification. Context management is embedded within the standard tool-call interface. When invoking $T_{\text{reconfig}}$, the inference agent produces the stage summary $H_{i}$ itself, so task execution and self-management share a single learnable pathway amenable to end-to-end optimization. This contrasts with ReSum (Wu et al., 2025b), which depends on an external model for periodic context compression.
@@ 3.1 Experimental Setup
Benchmarks (Sec. A.3). We evaluate ToolSelf on four benchmarks spanning diverse agentic capabilities: FRAMES (Krishna et al., 2024) and xbench (Chen et al., 2025) for deep research, GAIA (Mialon et al., 2023) for general AI assistant tasks, and SWE-bench Lite (Jimenez et al., 2023) for software engineering. For FRAMES (824 questions), we randomly partition 200 samples for testing and 624 for training. For GAIA, following WebSailor (Li et al., 2025a), we curate a subset of 103 samples that require only search and web browsing, denoted as GAIA(WS). Additionally, we incorporate 300 tasks from MathVista (Lu et al., 2023) to enhance multimodal capabilities.
Baselines (Sec. A.4). We compare against three categories of baselines: (1) Single-agent paradigms including Vanilla Agent equivalent to $\pi_{i}$ without $T_{\text{reconfig}}$, WebSailor (Li et al., 2025a) specialized for information retrieval, and SWE-Agent (Yang et al., 2024) designed for software engineering; (2) Single-agent extensions that introduce external mechanisms for partial improvement, including ReSum (Wu et al., 2025b) with periodic summarization and OAgents (Zhu et al., 2025) with periodic plan updates via an LLM planner; (3) Multi-agent workflows including OWL (Hu et al., 2025) with hierarchical architecture, Co-Sight (Zhang et al., 2025) with conflict-aware meta-verification, and SWE-Search (Antoniades et al., 2024) leveraging Monte Carlo tree search.
@@ 3.3 Configuration-Aware Two-stage Training
Table 2: Effect of Configuration-Aware Two-stage Training (CAT). Progressive performance gains through RFT cold-start initialization and KTO reinforcement learning refinement on Qwen3-8B. Vanilla Agent denotes a standard ReAct-style agent equivalent to $\pi_{i}$ without $T_{\text{reconfig}}$. GAIA(WS) refers to the WebSailor subset of GAIA containing tasks solvable via search and browsing only. WebSailor-7B denotes the specialized model from Li et al. (2025a). ToolSelf+RFT applies rejection sampling fine-tuning for cold-start initialization; ToolSelf+RFT+KTO further incorporates KTO reinforcement learning. Avg. denotes the average performance across all available benchmarks.
@@ 4 Related Work
Single-agents with static configurations. Single-agents, epitomized by ReAct (Yao et al., 2022), adopt minimally-prescriptive autonomous execution through iterative thought-action-observation cycles. Domain-specific variants include SWE-Agent (Yang et al., 2024) and OpenHands (Wang et al., 2025) for software engineering, and WebSailor (Li et al., 2025a), Search-o1 (Li et al., 2025b), and WebThinker-RL (Li et al., 2025c) for deep research. While offering flexibility across domains, these agents operate under static configurations fixed prior to execution. As analyzed in Sec. 1, this rigidity becomes a fundamental bottleneck in long-horizon settings: sub-goals cannot be decomposed adaptively, tool selection cannot respond to evolving needs, and context accumulates without principled management—ultimately manifesting as search space explosion and context poisoning.
@@
Multi-agent workflows with externally orchestrated configurations. Multi-agent workflows (Hu et al., 2025; Fourney et al., 2024) decompose tasks through human prior knowledge and assign specialized configurations to sub-agents via predefined orchestration. OWL (Hu et al., 2025) implements a hierarchical architecture where a Planner decomposes tasks, a Coordinator dispatches them to specialized Workers (e.g., Web Agent, Coding Agent), and re-plans upon failure. Magentic-One (Fourney et al., 2024) employs a similar orchestrator-workers pattern. Co-Sight (Zhang et al., 2025) enhances multi-agent reliability through conflict-aware meta-verification and structured fact management. WebWeaver (Li et al., 2025d) utilizes a Planner Agent followed by a Writer Agent for deep research. For software engineering, SWE-Search (Antoniades et al., 2024) leverages Monte Carlo tree search to systematically explore solution spaces. While effective in narrowly defined domains, these workflows rely on manually pre-specified configurations for each sub-agent, with orchestration logic hard-coded by human designers. This fundamentally limits their scalability and generalization: the configuration space is constrained by what designers anticipate, leaving them brittle when confronted with unfamiliar tasks or dynamically changing requirements (Qiu et al., 2025).
@@
Single-agent extensions with heuristic-triggered adjustment. Recent efforts introduce external mechanisms for partial configuration adjustment on single-agents. ReSum (Wu et al., 2025b) and MEM1 (Zhou et al., 2025) address context management through periodic summarization and learned compression strategies, respectively; OAgents (Zhu et al., 2025) employs an external LLM planner to periodically update the agent’s plan during execution; AdaPlanner (Sun et al., 2023) adapts plans based on environmental feedback. However, these approaches share two critical limitations: (1) they target isolated symptoms rather than providing unified configuration management, and (2) their triggering logic is governed by rigid heuristic thresholds—such as step counts or context length—which decouple configuration updates from task semantics and execution dynamics, thereby precluding end-to-end, data-driven optimization.

### v2

@@ Abstract
Agentic systems powered by Large Language Models (LLMs) have demonstrated remarkable potential in tackling complex, long-horizon tasks. However, their efficacy is fundamentally constrained by static configurations governing agent behaviors, which are fixed prior to execution and fail to adapt to evolving task dynamics. Existing approaches, relying on manual orchestration or heuristic-based patches, often struggle with poor generalization and fragmented optimization. To transcend these limitations, we propose ToolSelf, a novel paradigm enabling tool-driven runtime self-reconfiguration. By abstracting configuration updates as a callable tool, ToolSelf unifies task execution and self-adjustment into a single action space, achieving a phase transition from external rules to intrinsic parameters. Agents can thereby autonomously update their sub-goals and context based on task progression, and correspondingly adapt their strategy and toolbox, transforming from passive executors into dual managers of both task and itself. We further devise Configuration-Aware Two-stage Training (CAT), combining rejection sampling fine-tuning with trajectory-level reinforcement learning to internalize this meta-capability. Extensive experiments across diverse benchmarks demonstrate that ToolSelf rivals specialized workflows while generalizing to novel tasks, achieving a 24.1% average performance gain and illuminating a path toward truly self-adaptive agents.
@@ 1 Introduction
Agentic systems refer to large language model–based systems that go beyond passive text generation to autonomously plan, reason, and act in interactive environments through multi-step decision making (kolt2025governing, ). By tightly coupling internal reasoning with external actions—such as tool invocation, environment interaction, and intermediate state updates—agentic systems enable models to tackle complex, long-horizon tasks that cannot be addressed through single-shot inference (yao2022react, ). This paradigm has become increasingly important as real-world applications, including deep research (wu2025agentic, ), software engineering (yang2024swe, ), and system control (wang2024appbench, ; qin2025ui, ), demand sustained goal pursuit, adaptive planning, and iterative integration of information over time.
@@
Central to the behavior of an agentic system is its configuration, which defines the internal control variables governing how the agent reasons, acts, and interacts with its environment during execution. Configuration typically encompasses the agent’s execution strategy (e.g., persona or reasoning style) (chen2023autoagents, ), the available toolbox of actions (lu2025octotools, ), and the task-relevant knowledge or context maintained across steps (zhou2025mem1, ). These elements implicitly shape the agent’s action space, information flow, and decision-making dynamics, and thus play a decisive role in long-horizon performance (guo2024large, ). However, most existing agentic systems rely on a static configuration that is fixed prior to execution and remains unchanged throughout the task. While adequate for short or deterministic questions, this rigidity becomes a fundamental bottleneck in long-horizon settings, where sub-goals evolve, information requirements shift, and tool needs change over time (wolflein2025llm, ). Consequently, configuration is not merely an implementation detail, but a first-class component of agentic systems that fundamentally determines their autonomy, efficiency, and capacity for sustained reasoning.
@@
Prior work (Appendix 4) has explored two main directions to tackle these challenges. Multi-agent workflows (hu2025owl, ; fourney2024magentic, ) rely on manually specified priors to decompose tasks in advance and orchestrate sub-agents with specialized configurations for collaborative execution. While effective in narrowly defined domains, the heavy reliance on human-designed decomposition, prompts, and orchestration logic severely limits their scalability, transferability, and generalization to unseen tasks. In contrast, single-agent extensions introduce external mechanisms for periodic configuration adjustment (sun2023adaplanner, ; zhu2025oagents, ) or interaction history compression (wu2025resum, ; zhou2025mem1, ). However, these solutions operate as fragmented patches targeting isolated symptoms rather than addressing the underlying cause. More critically, their triggering logic is typically governed by rigid heuristic thresholds—such as step counts or context length—which decouple configuration updates from task semantics and execution dynamics, thereby precluding end-to-end, data-driven optimization.
@@
To this end, we propose ToolSelf (Tool-driven Self-reconfiguration), a novel paradigm that unifies task execution and self-configuration into a single action space by abstracting configuration updates as a callable tool, thereby effecting a phase transition from externally imposed rules to intrinsically learned control parameters. As illustrated in Figure 1, ToolSelf departs from conventional static architectures by equipping the inference agent with a reconfiguration tool, enabling it to autonomously and dynamically update its sub-goals (sun2023adaplanner, ) and context in response to task progression and environmental feedback, while adaptively configuring appropriate execution strategies and toolsets for each task stage. Sub-goal updates provide macro-level guidance that mitigates blind exploration in exponentially growing search spaces (zhou2023language, ), while proactive context management ensures refined and efficient information processing (wu2025resum, ; zhou2025mem1, ). Together, these mechanisms elevate agents from passive executors to dual managers of both task execution and self-configuration.
To enable the data-driven acquisition of self-reconfiguration capabilities, we further devise Configuration-Aware Two-stage Training (CAT). Both the inference agent and the reconfiguration engine share a single base model, yielding a compact and end-to-end trainable architecture. CAT optimizes both modules through two stages: Stage I employs rejection sampling fine-tuning (RFT) (li2025websailor, ) to initialize both modules using teacher-generated successful trajectories, as effective self-reconfiguration is difficult to explore from scratch. Stage II applies KTO-based reinforcement learning (ethayarajh2024kto, ) to further optimize reconfiguration decisions. Since self-reconfiguration is a meta-capability across execution stages whose quality only manifests through downstream task performance, we adopt trajectory-level credit assignment—propagating task success/failure to all reconfiguration decisions within the trajectory—for which KTO’s binary feedback is naturally well aligned.
@@ 2.1 Preliminaries: Pre-Fixed Configuration Paradigm
Formally, the behavior policy of a ReAct-style (yao2022react, ) single-agent can be described as a conditional distribution:
@@ 2.3 Configuration-Aware Two-stage Training
To internalize the meta-capability of self-reconfiguration, we propose Configuration-Aware Two-stage Training (CAT) to jointly train both the inference agent $\pi_{i}$ and the reconfiguration engine $\mu$. CAT consists of Rejection Sampling Fine-tuning (RFT) (li2025websailor, ) for cold-start initialization and Kahneman-Tversky Optimization (KTO) (ethayarajh2024kto, ) for reinforcement learning refinement. Both $\pi_{i}$ and $\mu$ are fine-tuned via Low-Rank Adaptation (LoRA) (hu2022lora, ) with independent adapters. Let $\theta_{0}$ denote the frozen base model parameters, and let $\Delta\theta_{\pi}$ and $\Delta\theta_{\mu}$ denote the LoRA parameters for the inference agent and reconfiguration engine, respectively. The effective parameters of the two modules are thus $\theta_{\pi}=\theta_{0}+\Delta\theta_{\pi}$ and $\theta_{\mu}=\theta_{0}+\Delta\theta_{\mu}$. Crucially, $\pi_{i}$ and $\mu$ share a single base model; LoRA adapters merely specialize it, yielding a compact and end-to-end trainable architecture.
Stage I: Cold-Start Initialization. To provide strong priors for configuration decisions, we bootstrap from teacher demonstrations (li2025websailor, ). We generate trajectories using a teacher model on the task dataset and filter successful trajectories via rejection sampling. For task $Q^{(j)}$, let $\mathcal{T}_{\text{teacher}}^{(j)}$ denote the successful trajectory obtained through rejection sampling, and define a binary success label $y(\mathcal{T}_{\text{teacher}}^{(j)})\in\{0,1\}$, where 1 indicates successful task completion and 0 indicates failure. The filtered dataset is defined as:
@@ 2.4 Design Advantages
Autonomy. The inference agent retains full authority over macro-level task management. By formulating reconfiguration requests $r_{i}$—including proposed sub-goals $q_{i+1}^{\text{prop}}$—the agent itself drives task decomposition and progress tracking; the reconfiguration engine $\mu$ merely acts as a “compiler” translating these requests into concrete configurations. This contrasts with OWL (hu2025owl, ) and OAgents (zhu2025oagents, ), which delegate planning to a separate high-level module, incurring additional coordination overhead.
Unification. Context management is embedded within the standard tool-call interface. When invoking $T_{\text{reconfig}}$, the inference agent produces the stage summary $H_{i}$ itself, so task execution and self-management share a single learnable pathway amenable to end-to-end optimization. This contrasts with ReSum (wu2025resum, ), which depends on an external model for periodic context compression.
@@ 3.1 Experimental Setup
Benchmarks (Sec. A.3). We evaluate ToolSelf on four benchmarks spanning diverse agentic capabilities: FRAMES (krishna2024fact, ) and xbench (chen2025xbench, ) for deep research, GAIA (mialon2023gaia, ) for general AI assistant tasks, and SWE-bench Lite (jimenez2023swe, ) for software engineering. For FRAMES (824 questions), we randomly partition 200 samples for testing and 624 for training. For GAIA, following WebSailor (li2025websailor, ), we curate a subset of 103 samples that require only search and web browsing, denoted as GAIA(WS). Additionally, we incorporate 300 tasks from MathVista (lu2023mathvista, ) to enhance multimodal capabilities.
Baselines (Sec. A.4). We compare against three categories of baselines: (1) Single-agent paradigms including Vanilla Agent equivalent to $\pi_{i}$ without $T_{\text{reconfig}}$, WebSailor (li2025websailor, ) specialized for information retrieval, and SWE-Agent (yang2024swe, ) designed for software engineering; (2) Single-agent extensions that introduce external mechanisms for partial improvement, including ReSum (wu2025resum, ) with periodic summarization and OAgents (zhu2025oagents, ) with periodic plan updates via an LLM planner; (3) Multi-agent workflows including OWL (hu2025owl, ) with hierarchical architecture, Co-Sight (zhang2025cosight, ) with conflict-aware meta-verification, and SWE-Search (antoniades2024swe, ) leveraging Monte Carlo tree search.
@@ 3.3 Configuration-Aware Two-stage Training
Table 2: Effect of Configuration-Aware Two-stage Training (CAT). Progressive performance gains through RFT cold-start initialization and KTO reinforcement learning refinement on Qwen3-8B. Vanilla Agent denotes a standard ReAct-style agent equivalent to $\pi_{i}$ without $T_{\text{reconfig}}$. GAIA(WS) refers to the WebSailor subset of GAIA containing tasks solvable via search and browsing only. WebSailor-7B denotes the specialized model from li2025websailor . ToolSelf+RFT applies rejection sampling fine-tuning for cold-start initialization; ToolSelf+RFT+KTO further incorporates KTO reinforcement learning. Avg. denotes the average performance across all available benchmarks.
@@ 4 Related Work
Single-agents with static configurations. Single-agents, epitomized by ReAct (yao2022react, ), adopt minimally-prescriptive autonomous execution through iterative thought-action-observation cycles. Domain-specific variants include SWE-Agent (yang2024swe, ) and OpenHands (wang2025openhands, ) for software engineering, and WebSailor (li2025websailor, ), Search-o1 (li2025search, ), and WebThinker-RL (li2025webthinker, ) for deep research. While offering flexibility across domains, these agents operate under static configurations fixed prior to execution. As analyzed in Sec. 1, this rigidity becomes a fundamental bottleneck in long-horizon settings: sub-goals cannot be decomposed adaptively, tool selection cannot respond to evolving needs, and context accumulates without principled management—ultimately manifesting as search space explosion and context poisoning.
@@
Multi-agent workflows with externally orchestrated configurations. Multi-agent workflows (hu2025owl, ; fourney2024magentic, ) decompose tasks through human prior knowledge and assign specialized configurations to sub-agents via predefined orchestration. OWL (hu2025owl, ) implements a hierarchical architecture where a Planner decomposes tasks, a Coordinator dispatches them to specialized Workers (e.g., Web Agent, Coding Agent), and re-plans upon failure. Magentic-One (fourney2024magentic, ) employs a similar orchestrator-workers pattern. Co-Sight (zhang2025cosight, ) enhances multi-agent reliability through conflict-aware meta-verification and structured fact management. WebWeaver (li2025webweaver, ) utilizes a Planner Agent followed by a Writer Agent for deep research. For software engineering, SWE-Search (antoniades2024swe, ) leverages Monte Carlo tree search to systematically explore solution spaces. While effective in narrowly defined domains, these workflows rely on manually pre-specified configurations for each sub-agent, with orchestration logic hard-coded by human designers. This fundamentally limits their scalability and generalization: the configuration space is constrained by what designers anticipate, leaving them brittle when confronted with unfamiliar tasks or dynamically changing requirements (qiu2025alita, ).
@@
Single-agent extensions with heuristic-triggered adjustment. Recent efforts introduce external mechanisms for partial configuration adjustment on single-agents. ReSum (wu2025resum, ) and MEM1 (zhou2025mem1, ) address context management through periodic summarization and learned compression strategies, respectively; OAgents (zhu2025oagents, ) employs an external LLM planner to periodically update the agent’s plan during execution; AdaPlanner (sun2023adaplanner, ) adapts plans based on environmental feedback. However, these approaches share two critical limitations: (1) they target isolated symptoms rather than providing unified configuration management, and (2) their triggering logic is governed by rigid heuristic thresholds—such as step counts or context length—which decouple configuration updates from task semantics and execution dynamics, thereby precluding end-to-end, data-driven optimization.